# Effective Community Detection Over Streaming Bipartite Networks

### v2

@@ 1. Introduction
Nowadays, the management of bipartite graphs that involve two distinct types of nodes (e.g., customers and products, visitors and check-in locations, users and clicked websites) has been extensively studied in many real applications such as online recommendation from user-product transaction graphs (Fan et al., 2022; Weng et al., 2022; Kou et al., 2023), location-based services with visitor-location check-in graphs (Hang et al., 2018; Zhou et al., 2021; Wang et al., 2017; Kim et al., 2022), and behavior analysis in user-webpage click graphs (Sun et al., 2009; Neria et al., 2017).
The community detection (CD) over such bipartite graphs has recently become increasingly important for bipartite graph analysis and mining (Barber, 2007; Bouguessa and Nouri, 2021; Yen and Larremore, 2020; Zhou and Amini, 2019), which uncovers hidden and structurally cohesive bipartite subgraphs (or groups). The CD problem over bipartite graphs has numerous useful real-world applications, including online product recommendations and marketing, user behavior analysis, and identifying malicious user groups.
Although many prior works (Wang et al., 2023; Abidi et al., 2023; Wang et al., 2021; Zhang et al., 2024; Wang et al., 2022; Zhang et al., 2021) studied the CD problem in static bipartite graphs, real-world bipartite graphs are usually dynamically changing over time (e.g., updates of purchase transactions, check-ins, or website browsing). Therefore, in this paper, we will consider the Community Detection over Streaming Bipartite Network (CD-SBN), upon graph updates (e.g., edge insertions/deletions or edge weight changes). Below, we provide a motivating example of our CD-SBN problem to detect malicious user groups (communities) on a user-website click bipartite graph.
@@
In addition to the example above, the CD-SBN problem has many other real applications. For example, in the customer-product transaction bipartite graph, we can utilize the CD-SBN results to identify communities of customers who have recently exhibited purchasing behaviors related to online advertising and marketing. Similarly, in the visitor-location check-in bipartite graph (e.g., from Yelp (Inc., 2024)), we can identify a group of visitors who frequently check in at some common points of interest (POIs) and provide them with group buying coupons/discounts (e.g., Groupon (Groupon, 2024)).
@@ 2.1. Streaming Bipartite Networks
Bipartite graphs have been widely used in many real applications, such as movie recommendations for users (Grujić, 2008), recommending places of interest (POIs) to tourists (Lang et al., 2022), or online product shopping recommendation for customers (Chen et al., 2022; Shan and Zhao, 2017; Yoon et al., 2020).
@@ 2.2. Basic Patterns in the Bipartite Graph
In this subsection, we give the definitions of basic patterns, wedge and butterfly (Gou and Zou, 2023), in the bipartite graph $G$.
@@
(Wedge, $\angle(u_{i},v,u_{j})$) Given two user vertices $u_{i},u_{j}$ $\in U(G)$ and an item vertex $v\in L(G)$ in a bipartite graph $G$, a wedge (Gou and Zou, 2023), $\angle(u_{i},v,u_{j})$, is a path $u_{i}\rightarrow v\rightarrow u_{j}$, where $u_{i}$, $v$, and $u_{j}$ are called start-, middle-, and end-vertex of the wedge, respectively. The weight of a wedge $\angle(u_{i},v,u_{j})$ is given by:
@@
(Butterfly (Gou and Zou, 2023), $\Join(u_{i},u_{j},v_{a},v_{b})$) Given two user vertices $u_{i},u_{j}\in U(G)$ and two item vertices $v_{a},v_{b}\in L(G)$ in a bipartite graph $G$, a butterfly, $\Join(u_{i},u_{j},v_{a},v_{b})$, is a complete bipartite subgraph containing wedges $\angle(u_{i},v_{a},u_{j})$ and $\angle(u_{i},v_{b},u_{j})$.
The butterfly structure (as given in Definition 2.4) has been widely used in many algorithms and applications (Sanei-Mehri et al., 2018; Xu et al., 2022b; Zhou et al., 2021) to detect communities with high cohesiveness in bipartite graphs. In Figure 2, $\Join(u_{2},u_{3},v_{2},v_{3})$ is an example of a butterfly.
@@ 4.5. Score Upper Bound Pruning
Based on the multinomial theorem (Bolton, 1968) with power $n=2$, we have:
@@ 10  $(M,key)$ = de-heap()($\mathcal{H}$);
Complexity Analysis: Let $PP^{(j)}$ be the pruning power (i.e., the percentage of node entries that can be pruned) at the $j$-th level of synopsis $Syn$, where $0\leq j\leq h$ (here $h$ is the height of the tree synopsis $Syn$). Denote $f$ as the average fanout of the synopsis nodes in $\mathcal{I}$. For the index traversal, the number of nodes that need to be accessed is given by $\sum_{j=1}^{h}f^{h-j+1}\cdot(1-{PP}^{(j)})$. Next, to obtain $g$ satisfying the constraints, we need to prune the item vertices without any query keyword. Therefore, it is necessary to recompute the support of the relevant edges and the user relationship score of any pair of adjacent user vertices. Let $\overline{n}$ be the number of pruned item vertices and $deg_{avg}$ be the average degree of the item vertices. The time cost of re-computation is $\overline{n}\cdot deg_{avg}^{2}$ (given in Section 5.2 and Section 5.4). Then, we can obtain the maximal $(k,r,\sigma)$-bitruss in $O(deg_{avg}^{2})$ with computed edge support, which is demonstrated in (Sarıyüce and Pinar, 2018). Thus, it takes $O(f^{h+1}\cdot(1-{PP}^{(0)})\cdot\overline{n}\cdot deg_{avg}^{2})$ to refine candidate seed communities, where ${PP}^{(0)}$ is the pruning power over $r$-hop subgraphs in leaf nodes. Therefore, the total time complexity of Algorithm 2 is given by $O\left(\sum_{j=1}^{h}f^{h-j+1}\cdot(1-{PP}^{(j)})+f^{h+1}\cdot(1-{PP}^{(0)})\cdot\overline{n}\cdot deg_{avg}^{2}\right)$.
@@ 6.3. Continuous CD-SBN Processing Algorithm
Complexity Analysis: Let $deg_{avg}$ be the average degree of the item vertices. For edge insertion and expiration, the time cost of user relationship score re-computation is $O(deg_{avg})$ and of edge support update is $O(deg_{avg}^{2})$. Since obtaining the maximal $(k,r,\sigma)$-bitruss with computed edge supports and user relationship scores (Sarıyüce and Pinar, 2018) requires $O(deg_{avg}^{2})$, it takes $O(|R_{t-1}|deg_{avg}^{2})$ to maintain the result set at t1t-1 and $O(deg_{avg}^{3})$ to detect potential communities. Therefore, the total time complexity of Algorithm 3 at each timestamp $t$ is given by $O\left(|R_{t-1}|deg_{avg}^{2}+deg_{avg}^{3}\right)$.
@@ 7.2. CD-SBN Performance Evaluation
CD-SBN Community Case Study (RQ3): To evaluate the utility of our CD-SBN results, we carry out a case study to compare our proposed $(k,r,\sigma)$-bitruss pattern with the biclique (Lin, 2018) (which is a variant of bipartite community structure) over BibSonomy (BS) graph data. The entire graph shown in Figure 6 is a $(4,2,3)$-bitruss containing 3 users (in blue) and 9 items (in blue or red), whereas the 3 users, 4 items (in blue) form a $3\times 4$ biclique. We can see that the $3\times 4$ biclique shares the same 3 users as the $(4,2,3)$-bitruss, but the $3\times 4$ biclique omits some items that are connected to users with large weights (e.g., 6). In contrast, our $(k,r,\sigma)$-bitruss captures these items, and can recommend them to users who are not connected to them (e.g., recommending items v_449272v\_449272, v_449273v\_449273, v_449277v\_449277, v_449275v\_449275, and $449276$ to user u_4954u\_4954). This confirms the utility of our proposed CD-SBN community semantics, which can recommend items to users who are potentially more interested.
@@ 8. Related Work
Community Search/Detection Over Unipartite Graphs: As fundamental operations in social network analysis, the community search (CS) and community detection (CD) have been extensively studied for the past few decades. In unipartite graphs, different community semantics have been proposed, including minimum degree (Cui et al., 2014), $k$-core (Sozio and Gionis, 2010), $k$-clique (Cui et al., 2013), and $(k,d)$-truss (Huang and Lakshmanan, 2017; Al-Baghdadi and Lian, 2020). In contrast, our work is conducted on bipartite graphs with two distinct types of nodes, and our proposed $(k,r,\sigma)$-bitruss considers the high connectivity of community structure, edge weights, and query keywords, which is more challenging.
@@
Dynamic Community Search/Detection Over Unipartite Graphs: Due to the importance of hidden trends in streaming data, CS/CD over large-scale dynamic graphs has been widely studied. Some recent studies aim to identify communities in dynamic social networks, e.g., the Triangle-connected $k$-Truss Community (Xu et al., 2022a) and the $(\theta,k)$-core (Tang et al., 2024, 2022). However, our CD-SBN problem aims to detect communities over streaming bipartite graphs, which is a more complicated task.
@@
Figure 6. A case study of $(k,r,\sigma)$-bitruss vs. biclique (Lin, 2018) (query keyword set $Q$ = {“Bookmarks”}).:
@@
Community Search/Detection Over Bipartite Graphs: Prior works on CD over static bipartite graphs include model-based  (Qing and Wang, 2023; Rohe et al., 2016) and modularity-based methods (Zhou et al., 2022; Bouguessa and Nouri, 2021), which require recalculations upon graph updates and are computationally expensive. Moreover, several dense-subgraph-based methods did not consider edge weights, including the size-bounded $(\alpha,\beta)$-core (Zhang et al., 2024; Wang et al., 2022), personalized $k$-wing (Abidi et al., 2023), and Pareto-optimal $(\alpha,\beta)$-community (Zhang et al., 2021). Other works focused on weighted bipartite graphs and considered significant $(\alpha,\beta)$-community (SC) (Wang et al., 2021, 2023) and $(\alpha,\beta)$-attributed weighted community ($(\alpha,\beta)$-AWC) (Li et al., 2022). However, the SC and the $(\alpha,\beta)$-AWC consider different community semantics from ours, so they cannot be directly utilized to solve our CD-SBN problem. Furthermore, Zhou et al.(Zhou et al., 2022) employed a learning-based clustering approach to identify communities on bipartite graphs; however, the resulting communities are sometimes unstable in terms of structure and size. The aforementioned works focused on static bipartite graphs, but do not apply to our CD-SBN problem on streaming bipartite graphs. Moreover, some prior works (Wang et al., 2023; Abidi et al., 2023) constructed a dynamically updated index to solve CS/CD on streaming bipartite graphs, requiring re-traversal of the entire index. In contrast, our CD-SBN approach incrementally maintains the result set upon edge updates with lower costs.

### v3

# Efficient Community Detection Over Streaming Bipartite Networks

@@ 1. Introduction
Nowadays, the management of bipartite graphs that involve two distinct types of nodes (e.g., customers and products, visitors and check-in locations, users and clicked websites) has been extensively studied in many real applications such as online recommendation from user-product transaction graphs (fan2022ModelingUserBehavior; weng2022distributed; kou2023ModelingSequentialCollaborative), location-based services with visitor-location check-in graphs (hang2018ExploringStudentCheckIn; zhou2021ButterflyCountingUncertain; wang2017SemanticAnnotationPlaces; kim2022abc), and behavior analysis in user-webpage click graphs (sun2009RankingbasedClusteringHeterogeneous; neria2017risk).
The community detection (CD) over such bipartite graphs has recently become increasingly important for bipartite graph analysis and mining (barber2007modularity; bouguessa2020BiNeTClusBipartiteNetwork; yen2020community; zhou2019analysis), which uncovers hidden and structurally cohesive bipartite subgraphs (or groups). The CD problem over bipartite graphs has numerous useful real-world applications, including online product recommendations and marketing, user behavior analysis, and identifying malicious user groups.
Although many prior works (wang2023DiscoveringSignificantCommunities; abidi2023SearchingPersonalizedkWing; wang2021EfficientEffectiveCommunity; zhang2024SizeboundedCommunitySearch; wang2022EfficientPersonalizedMaximum; zhang2021ParetooptimalCommunitySearch) studied the CD problem in static bipartite graphs, real-world bipartite graphs are usually dynamically changing over time (e.g., updates of purchase transactions, check-ins, or website browsing). Therefore, in this paper, we will consider the Community Detection over Streaming Bipartite Network (CD-SBN), upon graph updates (e.g., edge insertions/deletions or edge weight changes). Below, we provide a motivating example of our CD-SBN problem to detect malicious user groups (communities) on a user-website click bipartite graph.
@@
In addition to the example above, the CD-SBN problem has many other real applications. For example, in the customer-product transaction bipartite graph, we can utilize the CD-SBN results to identify communities of customers who have recently exhibited purchasing behaviors related to online advertising and marketing. Similarly, in the visitor-location check-in bipartite graph (e.g., from Yelp (yelp2024)), we can identify a group of visitors who frequently check in at some common points of interest (POIs) and provide them with group buying coupons/discounts (e.g., Groupon (groupon2024)).
@@ 2.1. Streaming Bipartite Networks
Bipartite graphs have been widely used in many real applications, such as movie recommendations for users (grujic2008MoviesRecommendationNetworks), recommending places of interest (POIs) to tourists (lang2022POIRecommendationBased), or online product shopping recommendation for customers (chen2022SocialBoostedRecommendation; shan2017AdaptiveKendallsCorrelationa; yoon2020ItemRecommendationPredicting).
@@ 2.2. Basic Patterns in the Bipartite Graph
In this subsection, we give the definitions of basic patterns, wedge and butterfly (gou2023SlidingWindowbasedApproximate), in the bipartite graph $G$.
@@
(Wedge, $\angle(u_{i},v,u_{j})$) Given two user vertices $u_{i},u_{j}$ $\in U(G)$ and an item vertex $v\in L(G)$ in a bipartite graph $G$, a wedge (gou2023SlidingWindowbasedApproximate), $\angle(u_{i},v,u_{j})$, is a path $u_{i}\rightarrow v\rightarrow u_{j}$, where $u_{i}$, $v$, and $u_{j}$ are called start-, middle-, and end-vertex of the wedge, respectively. The weight of a wedge $\angle(u_{i},v,u_{j})$ is given by:
@@
(Butterfly (gou2023SlidingWindowbasedApproximate), $\Join(u_{i},u_{j},v_{a},v_{b})$) Given two user vertices $u_{i},u_{j}\in U(G)$ and two item vertices $v_{a},v_{b}\in L(G)$ in a bipartite graph $G$, a butterfly, $\Join(u_{i},u_{j},v_{a},v_{b})$, is a complete bipartite subgraph containing wedges $\angle(u_{i},v_{a},u_{j})$ and $\angle(u_{i},v_{b},u_{j})$.
The butterfly structure (as given in Definition 2.4) has been widely used in many algorithms and applications (sanei-mehri2018ButterflyCountingBipartite; xu2022EfficientLoadbalancedButterfly; zhou2021ButterflyCountingUncertain) to detect communities with high cohesiveness in bipartite graphs. In Figure 2, $\Join(u_{2},u_{3},v_{2},v_{3})$ is an example of a butterfly.
@@ 4.5. Score Upper Bound Pruning
Based on the multinomial theorem (bolton1968MultinomialTheorem) with power $n=2$, we have:
@@ 10  $(M,key)$ = de-heap()($\mathcal{H}$);
Complexity Analysis: Let $PP^{(j)}$ be the pruning power (i.e., the percentage of node entries that can be pruned) at the $j$-th level of synopsis $Syn$, where $0\leq j\leq h$ (here $h$ is the height of the tree synopsis $Syn$). Denote $f$ as the average fanout of the synopsis nodes in $\mathcal{I}$. For the index traversal, the number of nodes that need to be accessed is given by $\sum_{j=1}^{h}f^{h-j+1}\cdot(1-{PP}^{(j)})$. Next, to obtain $g$ satisfying the constraints, we need to prune the item vertices without any query keyword. Therefore, it is necessary to recompute the support of the relevant edges and the user relationship score of any pair of adjacent user vertices. Let $\overline{n}$ be the number of pruned item vertices and $deg_{avg}$ be the average degree of the item vertices. The time cost of re-computation is $\overline{n}\cdot deg_{avg}^{2}$ (given in Section 5.2 and Section 5.4). Then, we can obtain the maximal $(k,r,\sigma)$-bitruss in $O(deg_{avg}^{2})$ with computed edge support, which is demonstrated in (sariyuce2018PeelingBipartiteNetworks). Thus, it takes $O(f^{h+1}\cdot(1-{PP}^{(0)})\cdot\overline{n}\cdot deg_{avg}^{2})$ to refine candidate seed communities, where ${PP}^{(0)}$ is the pruning power over $r$-hop subgraphs in leaf nodes. Therefore, the total time complexity of Algorithm 2 is given by $O\left(\sum_{j=1}^{h}f^{h-j+1}\cdot(1-{PP}^{(j)})+f^{h+1}\cdot(1-{PP}^{(0)})\cdot\overline{n}\cdot deg_{avg}^{2}\right)$.
@@ 6.3. Continuous CD-SBN Processing Algorithm
Complexity Analysis: Let $deg_{avg}$ be the average degree of the item vertices. For edge insertion and expiration, the time cost of user relationship score re-computation is $O(deg_{avg})$ and of edge support update is $O(deg_{avg}^{2})$. Since obtaining the maximal $(k,r,\sigma)$-bitruss with computed edge supports and user relationship scores (sariyuce2018PeelingBipartiteNetworks) requires $O(deg_{avg}^{2})$, it takes $O(|R_{t-1}|deg_{avg}^{2})$ to maintain the result set at t1t-1 and $O(deg_{avg}^{3})$ to detect potential communities. Therefore, the total time complexity of Algorithm 3 at each timestamp $t$ is given by $O\left(|R_{t-1}|deg_{avg}^{2}+deg_{avg}^{3}\right)$.
@@ 7.2. CD-SBN Performance Evaluation
CD-SBN Community Case Study (RQ3): To evaluate the utility of our CD-SBN results, we carry out a case study to compare our proposed $(k,r,\sigma)$-bitruss pattern with the biclique (lin2018TheParameterized) (which is a variant of bipartite community structure) over BibSonomy (BS) graph data. The entire graph shown in Figure 6 is a $(4,2,3)$-bitruss containing 3 users (in blue) and 9 items (in blue or red), whereas the 3 users, 4 items (in blue) form a $3\times 4$ biclique. We can see that the $3\times 4$ biclique shares the same 3 users as the $(4,2,3)$-bitruss, but the $3\times 4$ biclique omits some items that are connected to users with large weights (e.g., 6). In contrast, our $(k,r,\sigma)$-bitruss captures these items, and can recommend them to users who are not connected to them (e.g., recommending items v_449272v\_449272, v_449273v\_449273, v_449277v\_449277, v_449275v\_449275, and $449276$ to user u_4954u\_4954). This confirms the utility of our proposed CD-SBN community semantics, which can recommend items to users who are potentially more interested.
@@ 8. Related Work
Community Search/Detection Over Unipartite Graphs: As fundamental operations in social network analysis, the community search (CS) and community detection (CD) have been extensively studied for the past few decades. In unipartite graphs, different community semantics have been proposed, including minimum degree (cui2014LocalSearchCommunitiesa), $k$-core (sozio2010CommunitysearchProblemHowa), $k$-clique (cui2013OnlineSearchOverlapping), and $(k,d)$-truss (huang2017attribute; al2020topic). In contrast, our work is conducted on bipartite graphs with two distinct types of nodes, and our proposed $(k,r,\sigma)$-bitruss considers the high connectivity of community structure, edge weights, and query keywords, which is more challenging.
@@
Dynamic Community Search/Detection Over Unipartite Graphs: Due to the importance of hidden trends in streaming data, CS/CD over large-scale dynamic graphs has been widely studied. Some recent studies aim to identify communities in dynamic social networks, e.g., the Triangle-connected $k$-Truss Community (xu2022efficient) and the $(\theta,k)$-core (tang2024ReliabilityDrivenLocalCommunity; tang2022ReliableCommunitySearch). However, our CD-SBN problem aims to detect communities over streaming bipartite graphs, which is a more complicated task.
@@
Figure 6. A case study of $(k,r,\sigma)$-bitruss vs. biclique (lin2018TheParameterized) (query keyword set $Q$ = {“Bookmarks”}).:
@@
Community Search/Detection Over Bipartite Graphs: Prior works on CD over static bipartite graphs include model-based  (qing2023CommunityDetectionWeighted; rohe2016CoclusteringDirectedGraphs) and modularity-based methods (zhou2022EndtoendModularitybasedCommunity; bouguessa2020BiNeTClusBipartiteNetwork), which require recalculations upon graph updates and are computationally expensive. Moreover, several dense-subgraph-based methods did not consider edge weights, including the size-bounded $(\alpha,\beta)$-core (zhang2024SizeboundedCommunitySearch; wang2022EfficientPersonalizedMaximum), personalized $k$-wing (abidi2023SearchingPersonalizedkWing), and Pareto-optimal $(\alpha,\beta)$-community (zhang2021ParetooptimalCommunitySearch). Other works focused on weighted bipartite graphs and considered significant $(\alpha,\beta)$-community (SC) (wang2021EfficientEffectiveCommunity; wang2023DiscoveringSignificantCommunities) and $(\alpha,\beta)$-attributed weighted community ($(\alpha,\beta)$-AWC) (li2022AlphaBetaAWCSAlpha). However, the SC and the $(\alpha,\beta)$-AWC consider different community semantics from ours, so they cannot be directly utilized to solve our CD-SBN problem. Furthermore, Zhou et al.(zhou2022EndtoendModularitybasedCommunity) employed a learning-based clustering approach to identify communities on bipartite graphs; however, the resulting communities are sometimes unstable in terms of structure and size. The aforementioned works focused on static bipartite graphs, but do not apply to our CD-SBN problem on streaming bipartite graphs. Moreover, some prior works (wang2023DiscoveringSignificantCommunities; abidi2023SearchingPersonalizedkWing) constructed a dynamically updated index to solve CS/CD on streaming bipartite graphs, requiring re-traversal of the entire index. In contrast, our CD-SBN approach incrementally maintains the result set upon edge updates with lower costs.